Helen has two examinations in a month, one in History and one in French. She believes she has a probability of 0.78 of scoring an "A" grade in the History test. She decides that her performance on the two examinations are not independent. In particular, if she were to score an "A" grade on the History examination she believes her chance of an "A" in French would be 0.72. On the other hand, if she does not score an "A" on the History test, she thinks her chance of scoring an "A" on the French test is only 0.48. Provide answers to the following to two decimal places.

The result for the French examination comes out first. Helen learns she scored an "A" in the French test. Given this information, what is the probability she scored an "A" in the History test?
Answer #1

P(scored A on French)=P(Scored A on history and scored A on French)+P( Not Scored A on history and scored A on French)

=0.78*0.72+(1-0.78)*0.48 =0.6672

therefore P(Scored A on history given scored A on French)

=P(Scored A on history and scored A on French)/P(scored A on French)

=0.78*0.72/0.6672 =0.841727